Accelerator Update for May 26, 2008 - May 28, 2008
Monday May 26
The day shift began with Operations monitoring Tevatron (TeV) store 6165 (200.45E30) and the Recycler stash (52.6E10), with the Antiproton Source (Pbar) stacking (33.5E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, with the NuMI experiment off due to a LCW problem with their #1 horn, and with the SciBooNE / MiniBooNE (BNB) and Meson MTest experiments taking beam.
Operations had NuMI prepared for access by 8:28 AM.
At 9:38 AM, Operations reported that a MI RF station (MIRF3) was tripping off repeatedly. Operators contacted an expert who said that this was probably due to a specific event in the Time Line Generator (TLG) and told them how to adjust the station to work with this event.
A NuMI expert reported at 9:51 AM that the injector pump for the horn LCW system had failed. They won’t take any beam until the pump gets repaired, which might occur in a day or two.
At 11:09 AM, a Linac Klystron RF station (KRF6) tripped off. It reset.
At 11:26 AM, Operations reported a problem with beam falling out of MI at the start of the ramp. Operators investigated.
Operations reported at 11:52 AM that the Booster damper plot was flat during the MI problem. Operators investigated the dampers and the system began operating normally.
At 12:08 PM, KRF2 tripped off. It reset.
At 12:49 PM, Operations reported KRF4 & 7 trips were followed by Booster damper and MI beam permit trips.
At 1:02 PM, Operations reported that LRF3 had been causing problems for an hour.
At 1:05 PM, Operations reported that KRF5 was also down. It reset.
At 1:40 PM, KRF3 tripped off. It reset.
Operations reported at 4:02 PM that the Recycler had stashed antiprotons four times during the day shift.
At 5:53 PM, a TeV RF station (TRF7) tripped off. It will get reset after the store ends.
At 8:25 PM, Operations reported that they had lost control and readback for the MI RF vacuum section. Operations halted MI beam and investigated.
Operations put Pbar in standby at 9:04 PM.
At 9:42 PM, Operations reported that all of the big power supplies for F23 had tripped off. An expert suggested that a Yard Breaker had tripped. Operators contacted the Duty Electrician.
At 9:56 PM, Operations reported that they had replaced a RF CIA crate control card and reset all of the vacuum crates with no success. Operations called in an expert.
The Duty Electrician reported at 10:27 PM that the F23 Yard Breaker had tripped and he reset it.
Operations reported at 10:42 PM that the MI vacuum problems had been resolved and that Pbar had resumed stacking.
Operations reported at 10:48 PM that there seemed to have been a spike in the line voltage at the same time as the MI RF vacuum problems.
At 11:13 PM, KRF3 tripped off. It reset.
Operations reported at 11:56 PM that the Recycler had stashed five times during the evening shift.
Tuesday May 27
The midnight shift began with Operations monitoring store 6165 and the stash (325.8E10), with Pbar stacking (25.2E10), and with BNB taking beam.
Operations disabled the MTest beam at 12:10 AM.
Operations terminated store 6165 at 12:14 AM.
Operations turned TRF7 on at 12:23 AM.
Operations put the TeV into shot setup at 12:27 AM.
At 12:48 AM, KRF3 tripped off. It reset.
At 12:54 AM, Operations put the TeV shot setup on hold due to a problem with a TeV sector B48 vertical separator polarity switch. Operators investigated, checked cables, opened the door to the control rack, and then everything started working again. They carefully closed the door and slowly backed out of the room.
At 1:20 AM, LRF3 tripped off. It reset.
At 1:38 AM, a MI kicker (I:KPS5S) tripped off. Operators will reset the kicker later.
Operations established store 6166 at 2:25 AM, with an initial luminosity of 241.76E30.
At 3:42 AM, KRF6 tripped off. It reset.
Operations established beam to MTest at 6:11 AM.
Operations reported at 7:30 AM that they had run a compare program for the Recycler sequencer and discovered a number of devices out of tolerance. Recycler experts investigated and found a bad CAMAC crate card, but didn’t change it out. They will think about the best course of action.
A Booster expert began tuning the Booster at 7:52 AM.
At 7:59 AM, KRF6 tripped off. It reset.
Operations reported at 8:02 AM that the Recycler had stashed three times during the midnight shift.
At 8:08 AM, KRF1 tripped off. It reset.
At 8:55 AM, KRF2 tripped off. It reset.
At 9:53 AM, Operations reported a problem with the Recycler closure program.
At 9:55 AM, KRF2 tripped off. It reset.
A NuMI expert reported at 10:38 AM that they believed a ceramic isolator for the water suction line of horn #1 had failed. The solution to this problem would be to replace the horn, but this would take several weeks of work. Since this was primarily a cooling problem, they began working on several steps that might enable them to resume taking beam without replacing the horn.
A Booster expert began working on the GMPS at 10:50 AM.
At 10:59 AM, a power supply expert began working on a Lambertson power supply LCW leak.
At 11:38 AM, BRF11 tripped off. It reset.
A Linac expert reported at 11:51 AM that he had replaced the driver vacuum pump for LRF3.
Operations reported at 12:32 PM that certain Recycler sequencer commands were disabled and would remain disabled until experts replaced the 453 corrector card.
Operations reported at 1:30 PM that a NuMI expert was making up water for #1 horn LCW system.
A MI expert began a study period at 1:32 PM.
At 1:46 PM, the MI South Anode power supply tripped off. The power supply reset, but the VCB would not close. An expert investigated. Later, the VCB was closed.
At 3:21 PM, KRF2 tripped off. It reset.
Operations terminated store 6166 at 4:01 PM.
Operations reported at 4:02 PM that the Recycler had stashed five times during the day shift.
A TeV expert working on the TeV sector B48 separator reported at 4:15 PM that he found and repaired a poorly crimped wire connector. The B48 separator came back online.
Operations put the TeV in shot setup at 4:40 PM.
Operations began turning NuMI on at 4:44 PM. A quadrupole power supply (E:QF115) tripped off. Operators investigated and found smoke coming from the supply. They breakered off the supply and called in a power supply expert.
At 5:19 PM, a MI kicker (I:KPS5S) tripped off for the first of many times in about 17 minutes. It reset.
At 5:58 PM, KRF6 tripped off. It reset.
At 6:14 PM, the TeV suffered a quench in sectors DA and DB. Operators investigated.
At 6:20 PM, LRF4 tripped off for the first of two times six minutes. It reset.
Operations reported at 6:32 PM that there were some beam position problems at TeV sector D0 just prior to the quench.
At 7:19 PM, the F23 yard breaker tripped again. The Duty Electrician investigated and confirmed the breaker had tripped.
Cryo system experts began work on the TeV sector F3 wet engine at 7:45 PM.
A TeV expert came in at 8:20 PM to help investigate the quench. Later, he reported that the problem might have been due to an interaction between the protons and antiprotons.
The Duty Electrician reported at 9 PM that he had reset the F23 yard breaker.
Operations reported that E:QF115 had been repaired. NuMI resumed taking beam at 9:04 PM.
Pbar resumed stacking at 9:07 PM.
The TeV sector A1 cryo system resumed normal operation around 9:26 PM.
Operations prepared D0 and CDF for access at 10:04 PM.
At 11:08 PM, a power supply’s current (I:VTF11) drifted high and caused the P2 transfer line’s beam permit to trip off. Everything reset.
CDF and D0 came out of access at 11:31 PM.
Operations reported at midnight that the Recycler had stashed three times during the evening shift.
Wednesday May 28
The midnight shift began with Operations monitoring the stash (109.5E10), with Pbar stacking (20.4E10), and with NuMI, BNB, and MTest taking beam.
Operations disabled the MTest beam at 12:02 AM. This completed the CALICE T978 data taking for this portion of their experiment. They will return to Fermilab in July 2008.
At 12:27 AM, KRF2 tripped off. It reset.
Cryo system technicians reported at 1:04 AM that they had completed their F3 wet engine repairs.
At 2:12 AM, Operations reported that during the last stash, the F23 yard breaker tripped off and the transfer was lost. Operations contacted a power supply expert, the FESS Supervisors, and the High Voltage Coordinator.
Operations put Pbar into standby at 2:37 AM.
At 2:56 AM, MIRF18 tripped off. It reset.
Operations began ramping the TeV at 3:37 AM.
Operations put the TeV through a dry squeeze at 3:49 AM.
Operations reported at 4:38 AM that a power supply expert and the HV Coordinator replaced the F23 trip unit, but this did not fix the problem. The breaker needs to be replaced. FESS can not get an electrician in until around 6 AM.
At 6:21 AM, a Booster vacuum pump (IPRF17) tripped off. It reset.
At 7:17 AM, KRF6 tripped off. It reset.
The HV Coordinator reported that the F23 Yard Breaker replacement began at 7:19 AM.
Operations reported at 7:53 AM that the Recycler had stashed two times during the midnight shift.
At 7:59 AM, KRF2 tripped off. It reset.
The Plans for Wednesday
The plans for today are to replace the F23 Yard Breaker; to allow TeV experts to work on orbits. When the orbit studies end, Operations will establish a store.
Comments and Suggestions
|last modified 05/28/2008 email Fermilab|
|Security, Privacy, Legal|